OpenStudy (anonymous):

Domain and range determine the domain and range Y=2× y=8×+1

OpenStudy (gibbs):

x=y−18

OpenStudy (gibbs):

x = y -1/18

OpenStudy (anonymous):

U kn its 2 diff prob y=2× and y=8×+1

OpenStudy (anonymous):

Not all together